OpenStudy (lgbasallote):

i just need a demonstration so steps (no answer) would be required obtain the particular solution satisfying the initial condition indicated \[2xyy' = 1 + y^2\] when \(x = 2, y = 3\)

OpenStudy (turingtest):

\[2xyy'=1+y^2\]\[2xy\frac{dy}{dx}=1+y^2\]separate the variables. That is to say treat the derivative in Leibniz notation as a fraction and split it up. Get all the y's on one side and the x's on the other\[{ydy\over1+y^2}={2dx\over x}\]now integrate both sides...

OpenStudy (turingtest):

sorry the 2 is on the wrong side\[{2ydy\over1+y^2}={dx\over x}\]the left side requires a u-sub it makes no difference where you put the integration constant C; I'll put it on the RHS\[\ln

OpenStudy (lgbasallote):

\[\ln (1 + x^2) = \ln x + c\] that right?

OpenStudy (turingtest):

exactly :) don't forget the integration constant! this is where we finally use it

OpenStudy (turingtest):

oh but that should have a y on the LHS typo I'm sure

OpenStudy (lgbasallote):

lol yeah =)) so \[\ln (\frac{1 + y^2}{x}) = c\] substitute right?

OpenStudy (turingtest):

you could do it that way...

OpenStudy (lgbasallote):

there's another?

OpenStudy (turingtest):

but I wouldn't just in case x=0 is the initial condition ;)

OpenStudy (turingtest):

exponentiate both sides

OpenStudy (lgbasallote):

but x = 2 o.O

OpenStudy (turingtest):

yes, but it may not always be

OpenStudy (lgbasallote):

hmmm so then how do you exponentiate?

OpenStudy (turingtest):

\[\ln (1 + y^2) = \ln x + C\]raise e to the power of each side to undo the logs

OpenStudy (lgbasallote):

so \[1 + y^2 = e^{\ln x + C}??\]

OpenStudy (turingtest):

yes, and don't forget that\[x^{a+b}=x^a\cdot x^b\]apply that to the RHS

OpenStudy (lgbasallote):

\[1 + y^2 = x + e^c\]

OpenStudy (turingtest):

careful...

OpenStudy (turingtest):

\[1+y^2=e^c x\]

OpenStudy (lgbasallote):

oh..oh yeah

OpenStudy (turingtest):

but \(e^c\) is a constant, so we can just call it C again :P\[1+y^2=Cx\]

OpenStudy (lgbasallote):

so \[e^c = \frac{1+y^2}{x}\] now substitute?

OpenStudy (lgbasallote):

so \[C_1 = \frac{1+y^2}{x}\]

OpenStudy (turingtest):

I recommend keeping x and y on separate sides; again if the initial condition was x=0 you'd be stuck here with that move

OpenStudy (turingtest):

I'm trying to teach you a way that will always work, we could make exceptions here but that is a bad habit

OpenStudy (lgbasallote):

but i need to get C to get particular solution right?

OpenStudy (turingtest):

yes, but solving for c is usually quite easy by starting from the original equation\[1+y^2=Cx\implies 1+9=2C\implies C=5\]if x=0 we would have gotten 10=0, which would show no such C could exist I suppose solving for C right away like you did suggests the same thing, but I think it is more clear the way I showed you it's up to you

OpenStudy (lgbasallote):

hmm well okay...since c = 5...what does that make the particular solution?

OpenStudy (turingtest):

what was the general solution? put a 5 where the C was that's the particular solution

OpenStudy (lgbasallote):

general solution is \[\ln (1+y^2) = \ln x + c\] right?

OpenStudy (turingtest):

if you want it in that form you have to put the absolute value bars or you will change the domain

OpenStudy (lgbasallote):

\[\ln |1 + y^2| = \ln |x| + c\] is the general solution?

OpenStudy (turingtest):

yes, but it is equally valid exponentiated, so I would just write it as\[1+y^2=Cx\]

OpenStudy (turingtest):

again, whatever floats your boat (or whatever your teacher prefers)

OpenStudy (lgbasallote):

he doesnt prefer anything lol so basically i find the general solution..then from that solution i find C then substitute C into the general solution..that right?

OpenStudy (turingtest):

exactly :)

OpenStudy (lgbasallote):

let's say x = 2 and y = 3 are not given..what would the answer be then? would it just be the general solution?

OpenStudy (turingtest):

yes, that would just leave us with the integration constant C undetermined, which, if you continue to study mathematics, corresponds to a "family" of function dependent on the possible values of C

OpenStudy (turingtest):

...which is the general solution so the answer to your Q is "yes"

OpenStudy (lgbasallote):

good thanks :DD

OpenStudy (turingtest):

welcome you picked it up easily as usual I'm out, later!